# Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v1

@@ Abstract
We propose denoising diffusion variational inference (DDVI), an approximate inference algorithm for latent variable models which relies on diffusion models as expressive variational posteriors. Our method augments variational posteriors with auxiliary latents, which yields an expressive class of models that perform diffusion in latent space by reversing a user-specified noising process. We fit these models by optimizing a novel lower bound on the marginal likelihood inspired by the wake-sleep algorithm. Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. When applied to deep latent variable models, our method yields the denoising diffusion VAE (DD-VAE) algorithm. We use this algorithm on a motivating task in biology—inferring latent ancestry from human genomes—outperforming strong baselines on the Thousand Genomes dataset.
@@ 1 Introduction
Latent variable methods often rely on variational inference to fit an approximate model of the posterior distribution (Vahdat & Kautz, 2020; Maaløe et al., 2016). The expressivity of this model has a significant impact on the performance of variational inference (Kingma et al., 2016), which motivates research that leverages modern generative models—including normalizing flows (Rezende & Mohamed, 2015) and generative adversarial networks (Goodfellow et al., 2014; Makhzani et al., 2015)—to represent expressive approximate posteriors.
@@
This work seeks to improve variational inference via expressive posteriors based on diffusion models (Ho et al., 2020; Song et al., 2020). Diffusion methods are defined via a noising process, which maps data into Gaussian noise; a diffusion model generates data by reversing this noising process, which yields high-quality samples and accurate density estimates (Kingma et al., 2021). Here, we argue for using diffusion models in latent space, where we gradually map a simple (e.g., Gaussian) latent representation of the data into one that is more complex via an iterative diffusion-like procedure. This procedure yields an expressive approximate posterior trained with a denoising objective that does not involve adversarial training (Makhzani et al., 2015) or constrained invertible architectures (Kingma et al., 2016).
@@
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm that introduces auxiliary latent variables into the approximate posterior via a user-specified noising process. This process transforms the latent variable we seek to model into a simple (e.g., Gaussian) auxiliary latent; during inference, we fit the approximate posterior by reversing this noising process. Our learning objective is a variational lower bound inspired by the wake-sleep algorithm (Hinton et al., 1995) that can be interpreted as a form of regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
@@
Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. When applied to deep latent variable models, our method yields the denoising diffusion VAE (DD-VAE) algorithm. We evaluate DD-VAEs on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on the Thousand Genomes dataset (Siva, 2008) and learns a low-dimensional latent space that preserves semantically meaningful structure (Haghverdi et al., 2015).
@@
In summary, this work introduces denoising diffusion variational inference, an approximate inference algorithm that features three key components: auxiliary latent variables, a user-specified noising process over these variables, and a lower bound on the marginal likelihood inspired by wake-sleep. Our method can be used to fit deep latent variable models, which yields the DD-VAE algorithm. This algorithm is especially effective at dimensionality reduciton and representation learning, where it outperforms alternative methods based on adversarial training.
@@ 2 Background
$\log p_{\bm{\theta}}(\mathbf{x})\leq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}))$
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO (Kingma & Welling, 2013).
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods (Maaløe et al., 2016), MCMC-based methods (Salimans et al., 2015), normalizing flows (Rezende & Mohamed, 2015)—or improved learning objectives—e.g., adversarial or sample-based losses (Makhzani et al., 2015; Zhao et al., 2017; Si et al., 2022; 2023).
The wake-sleep algorithm (Hinton et al., 1995) optimizes an alternative objective
$\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))$
, in which the KL divergence term is reversed. The learning procedure for wake-sleep involves alternating between ”wake” phases where the recognition model is updated and ”sleep” phases where the generative model is refined.
@@
The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\bm{\theta}}$ with parameters ${\bm{\theta}}$ trained by maximizing the ELBO:
$\log p_{\bm{\theta}}(\mathbf{x}_{0})\geq\mathbb{E}_{q}\left[\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}D_{\mathrm{KL}}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t}))\right]-D_{\mathrm{KL}}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$
Visualization and Dimensionality Reduction
There exist two important types of dimensionality reduction methods, which (1) emphasize the preservation of pairwise distance structures among all data samples, e.g., PCA (Wold et al., 1987) and LDA (Balakrishnama & Ganapathiraju, 1998), and (2) local distances over global ones, e.g., t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018). Latent variable models (LVMs) (Kingma et al., 2016; Makhzani et al., 2015) are another class of techniques for dimensionality reduction. They represent high-dimensional data in terms of latent, or hidden, variables in a lower-dimensional space, effectively providing a compact representation of the data.
@@ 3 Variational Inference With Denoising Diffusion Models
We introduce denoising diffusion variational inference (DDVI), which enhances variational inference with diffusion-based methods and is motivated by challenges in data visualization and dimensionality reduction. Our approach consists in augmenting variational inference in a latent variable model $p(\mathbf{x},\mathbf{z})$ with auxiliary latents $\mathbf{y}\in\mathcal{Y}$ introduced via a user-specified noising process $r(\mathbf{y}|\mathbf{z})$. The $r(\mathbf{y}|\mathbf{z})$ transforms $\mathbf{z}$—which is the latent whose intractable posterior we seek to approximate—into $\mathbf{y}$, whose posterior will be easier to model. Examples of $r$ include forward diffusion processes, discrete noising processes (Austin et al., 2021), as well as custom regularizers (Section 3.3).
We then form an expressive posterior $q(\mathbf{z}|\mathbf{x})$ by fitting the reverse of the noising process $r(\mathbf{y}|\mathbf{z})$ as in a diffusion model. We define $q(\mathbf{z}|\mathbf{x})$ by sampling from a first model $q(\mathbf{y}|\mathbf{x})$—this is an easier task since we can choose $y$ to have a simple (e.g., Gaussian) posterior—and then by sampling from a denoising model $q(\mathbf{z}|\mathbf{x},\mathbf{y})$ that approximates the reverse process $r(\mathbf{z}|\mathbf{y})$. The model $q(\mathbf{z}|\mathbf{x},\mathbf{y})$ is parameterized and trained as a denoising diffusion model, and is thus a flexible posterior estimator.
When diffusion variational inference is used to fit a variational autoencoder, we refer to the resulting algorithm as VAE with diffusion encoders (DD-VAE). We define the full algorithm below.
3.1 DD-VAE: Variational Autoencoders With Diffusion Encoders
We seek to fit a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ with a potentially complex prior $p_{\bm{\theta}}(\mathbf{z})$. One perspective that can be used to define our approach consists in lifting $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ into an extended latent space by introducing auxiliary latents $\mathbf{y}\in\mathcal{Y}$ and applying variational inference in the extended space.
Specifically, we augment the model with the aforementioned user-specified noising process $r(\mathbf{y}|\mathbf{z})$. We use $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})=r(\mathbf{y}|\mathbf{z})p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ to denote the extended probability distribution. Note that the marginalizing out $\mathbf{y}$ in $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ yields the original model: hence fitting $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ in an extended probability space solves our original task.
@@ 3.1 DD-VAE: Variational Autoencoders With Diffusion Encoders
3.1.1 Denoising Diffusion Variational Inference (DDVI)
One possible way to apply variational inference to $p_{\bm{\theta}}(\mathbf{x},\mathbf{y},\mathbf{z})$ is to apply the ELBO twice to obtain:
$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{y})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y}))$ (1)
$\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x},\mathbf{y})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{y}))]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{y}))$ (2)
@@ 3.1.1 Denoising Diffusion Variational Inference (DDVI)
While Equation 2 is a valid learning objective, it does not yield a training procedure comparable to that of a diffusion model. Diffusion model training involves sampling noisy data from the forward process; here, we sample from the approximate reverse process (from $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$) and seek to match the true process $r(\mathbf{z}|\mathbf{y})$. In practice, we have found the learning signal from this procedure to be too weak to learn a good $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$ that reverses the noising process $r(\mathbf{y}|\mathbf{z})$, as illustrated in Table 7.
Instead, we adopt a learning objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ inspired by the wake-sleep algorithm:

### v3

# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

@@ Abstract
We propose denoising diffusion variational inference (DDVI), a black-box variational inference algorithm for latent variable models which relies on diffusion models as flexible approximate posteriors. Specifically, our method introduces an expressive class of diffusion-based variational posteriors that perform iterative refinement in latent space; we train these posteriors with a novel regularized evidence lower bound (ELBO) on the marginal likelihood inspired by the wake-sleep algorithm. Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. We find that DDVI improves inference and learning in deep latent variable models across common benchmarks as well as on a motivating task in biology—inferring latent ancestry from human genomes—where it outperforms strong baselines on the Thousand Genomes dataset.
@@ 1 Introduction
We are interested in amortized black-box variational inference problems of the form
$\log\;p_{\bm{\theta}}(\mathbf{x})\geq\max_{\bm{\phi}}\;\;\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log\;p_{\bm{\theta}}(\mathbf{x},\mathbf{z})-\log\;q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})\right]:=\max_{\bm{\phi}}\;\;\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}}),$ (1)
in which we approximate the marginal likelihood $\log\;p_{\bm{\theta}}(\mathbf{x})$ of a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$ with an evidence lower bound $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ that is a function of an approximate variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$. We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks, topic models [4], variational autoencoders (VAEs), and broad classes of $p_{\bm{\theta}}$ defined via modern probabilistic programming frameworks [9].
Maximizing $\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ over ${\bm{\phi}}$ yields a variational posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that approximates $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ as well as a tight bound on $\log\;p_{\bm{\theta}}(\mathbf{x})$ that serves as a learning objective for $p_{\bm{\theta}}$. The approximation gap $\log\;p_{\bm{\theta}}(\mathbf{x})-\max_{\bm{\phi}}\text{ELBO}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ equals precisely $\min_{\bm{\phi}}\text{KL}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$, which motivates the design of expressive classes of posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ that reduce this gap. Recent efforts leverage modern generative models—including normalizing flows [28, 16] and generative adversarial networks [8, 20]—as expressive model families for $q_{\bm{\phi}}$ that tighten the ELBO.
@@
This work seeks to further improve variational inference via expressive posteriors based on diffusion models [12, 37]. Diffusion methods have become the de-facto standard for high-quality image synthesis [30]. Here, we use diffusion in latent space to parameterize $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$. We train this distribution with a denoising diffusion-like objective that does not involve adversarial training [20] or constrained invertible normalizing flow architectures [16]. Samples from $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ are obtained via iterative refinement of $\mathbf{z}$, starting from a Gaussian distribution, and gradually forming one that is multi-modal and complex.
[…]
@@
Specifically, we propose denoising diffusion variational inference (DDVI), an approximate inference algorithm defined by a class of approximate posterior distribution based on diffusion as well as a learning objective inspired by the wake-sleep algorithm [11] that can be interpreted as a regularized variational inference. We also derive extensions of our method to semi-supervised learning and clustering.
@@
Our method is easy to implement (it fits a regularized extension of the ELBO), is compatible with black-box variational inference, and outperforms alternative classes of approximate posteriors based on normalizing flows or adversarial networks. We evaluate DDVI on synthetic benchmarks and on a real problem in biological data analysis—inferring human ancestry from genetic data. Our method outperforms strong baselines on the Thousand Genomes dataset [35] and learns a low-dimensional latent space that preserves biologically meaningful structure [10].
@@
In summary, this work introduces denoising diffusion variational inference, an approximate inference algorithm defined by two components: a class of approximate posteriors $q(\mathbf{z}|\mathbf{x})$ parameterized by diffusion, and a lower bound on the marginal likelihood inspired by wake-sleep. We complement DDVI with extensions to semi-supervised learning and clustering. Our method is especially suited for probabilistic programming, representation learning, and dimensionality reduction, where it outperforms alternative methods based on normalizing flows and adversarial training.
@@ 2 Background
$\log p_{\bm{\theta}}(\mathbf{x})\geq\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z})),$ (2)
which serves as a tractable surrogate for the marginal log-likelihood (MLL). The gap between the MLL and the ELBO equals precisely $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$—thus, a more expressive $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may better fit the true posterior and induce a tighter ELBO [14].
Expressive variational posteriors can be formed by choosing more expressive model families—including auxiliary variable methods [19], MCMC-based methods [31], normalizing flows [28]—or improved learning objectives—e.g., adversarial or sample-based losses [20, 48, 33, 34].
The wake-sleep algorithm [11] optimizes an alternative objective
$\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})),$ (3)
in which the KL divergence term is reversed. The learning procedure for wake-sleep involves alternating between "wake" phases where the recognition model is updated and "sleep" phases where the generative model is refined.
@@
The true reverse of the process $q$ cannot be expressed in closed form; as such, we parameterize $p_{\bm{\theta}}$ with ${\bm{\theta}}$ trained by maximizing the ELBO:
$\log p_{\bm{\theta}}(\mathbf{x}_{0})\geq\mathbb{E}_{q}\Bigr{[}\log p_{\bm{\theta}}(\mathbf{x}_{0}|\mathbf{x}_{1})-\sum_{t=2}^{T}D_{\mathrm{KL}}(q_{t}||p_{t})\Bigr{]}-D_{\mathrm{KL}}(q(\mathbf{x}_{T}|\mathbf{x}_{0})||p(\mathbf{x}_{T}))$ (4)
where $q_{t},p_{t}$ denote the distributions $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ and $p_{\bm{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, respectively.
@@ 3 Variational Inference With Denoising Diffusion Models
We introduce denoising diffusion variational inference (DDVI), which improves variational inference with diffusion-based techniques. The goal of DDVI is to fit a latent variable model $p_{\bm{\theta}}(\mathbf{x},\mathbf{z})$. We assume that $p_{\bm{\theta}}$ factorizes as $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})p_{\bm{\theta}}(\mathbf{z})$ and admits efficient sampling: examples of such $p_{\bm{\theta}}$ include Bayesian networks and variational autoencoders (VAEs) [14]. Our approach is comprised of two key components: a family of approximate posteriors $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ based on diffusion and a learning objective that can be seen as a regularized ELBO. The $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ iteratively refines latents $\mathbf{z}$, starting from a Gaussian distribution. The learning objective trains $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to reverse a used-specified forward diffusion process.
3.1 Diffusion-Based Variational Posteriors
@@
We choose the latent $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ to be a vector of $T$ variables that represent progressively simplified versions of $\mathbf{z}$, with $\mathbf{y}_{T}$ corresponding to a simple distribution (e.g., a Gaussian). The model $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$ transforms $\mathbf{y}_{T}$ into $\mathbf{z}$ via iterative refinement. To sample from $q_{\bm{\phi}}$, we first sample $\mathbf{y}_{T}$—this is an easier task since we can define $\mathbf{y}_{T}$ to have a simple (e.g., Gaussian) distribution—and then by sampling from the denoising model $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})\prod_{t=1}^{T-1}q_{\bm{\phi}}(\mathbf{y}_{t}|\mathbf{y}_{t+1},\mathbf{x})$.
We define the relationship between $\mathbf{y}$ and $\mathbf{z}$ via a forward diffusion process $r(\mathbf{y}|\mathbf{z})=r(\mathbf{y}_{1:T}|\mathbf{z})=r(\mathbf{y}_{1}|\mathbf{z})\prod_{t=1}^{T-1}r(\mathbf{y}_{t+1}|\mathbf{y}_{t})$, which transforms $\mathbf{z}$—the latent whose intractable posterior we seek to approximate—into $\mathbf{y}_{T}$, whose posterior is easier to model. Examples of $r$ include Gaussian forward diffusion processes and discrete noising processes [2]. The model $q_{\bm{\phi}}$ is trained to approximately reverse this forward diffusion process.
3.2 The Wake-Sleep Regularized ELBO
The standard approach to fit auxiliary-variable generative models [19] is to apply the ELBO twice:
$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$ (5)
$\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))-\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$ (6)
$=~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$ (7)
In Equation 5, we applied the ELBO over $\mathbf{z}$, and in Equation 6 we applied the ELBO again over the latent $\mathbf{y}$ of $q$. Notice that the gap between the ELBO and $\log p_{\bm{\theta}}(\mathbf{x})$ is $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$. Thus, if we correctly match $q$ and $r$, we will achieve a tight bound.
Wake-Sleep Regularization
@@ 3.2 The Wake-Sleep Regularized ELBO
Notice that optimizing $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{z},\mathbf{x})p(\mathbf{z}))$ in Equation 7 involves sampling from the approximate reverse process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$. We have found the standard ELBO to insufficient to learn a good $q_{\bm{\phi}}$ that reverses the noising process $r$, as illustrated in Table 8.
Instead, our strategy is to introduce a modified ELBO that yields diffusion-like training. This objective is the ELBO in Equation 7 augmented with an additional regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$.